OpenStudy (anonymous):

3/4 + x = 7/8 solve the following equations....I'm double checking my sons work....have a few problems....thanks

OpenStudy (anonymous):

3/4 + x = 7/8 Move 3/4 to the right side: x = 7/8 - 3/4 Make the denominators equal: x = 7/8 - 2(3/4) x = 7/8 - 6/8 7 - 6 is 1 so: x = 1/8